# Lifting the Veil on the Large Language Model Supply Chain: Composition, Risks, and Mitigations

### v2

@@ 5.3. Mitigation to Model Risks
Model Watermark. Model watermark embeds traceable information into the text generated by LLMs to verify the model’s source and authenticity without degrading output quality. Techniques such as logits modification (Kirchenbauer et al., 2023), entropy-based embedding (Lee et al., 2023), multi-bit payload (Wang et al., 2023), and token sampling (Kuditipudi et al., 2023) enable efficient watermarking by embedding verifiable identifiers that protect against model-stealing, safeguarding intellectual property and ensuring model integrity. This measure can mitigate the risk PR3.
@@
Federated Learning. Federated Learning is a decentralized approach where multiple clients collaboratively train a shared model without transferring their individual data to a central server. It prevents the aggregation of malicious updates by identifying and excluding anomalies in the distribution using unsupervised or supervised machine learning techniques  (Li et al., 2020; Zhang et al., 2022). By leveraging three-stage defenses (pre-aggregation defense (Pre-AD), in-aggregation defense (In-AD), and post-aggregation defense (Post-AD)), federated learning ensures that the aggregated model remains secure and resistant to backdoor attacks or malicious data manipulation, while still benefiting from decentralized data training. This measure can mitigate the risk SR3.
5.4. Mitigation to Prompt Risks
Prompt Sanitization. Prompt sanitization filters, formats, or regularized prompts are provided to LLMs to ensure that the model does not misinterpret their original intent, thereby preventing abnormal content or malicious behavior. (Suo, 2024; Robey et al., 2023) have been shown to effectively prevent jailbreak in LLMs. Suo et al. (Suo, 2024) applied to sign to sensitive elements (e.g.,’delete file’) in the user’s prompt. Robey et al. (Robey et al., 2023) randomly perturb characters in a given input prompt (e.g., inserting random characters). Chen et al .(Chen et al., 2024) separates prompts and user data into two independent channels to prevent prompt injection. This measure can mitigate the risk SR3.
@@ 5.4. Mitigation to Prompt Risks
Model Structure Optimization. Model structure optimization aims to enhance the robustness of LLMs’ inference by modifying their structure to defend against malicious prompts. Chen et al (Chen et al., 2022). Employ homomorphic encryption to enable privacy-preserving inference for Transformer-based models. Li et al. (Li et al., 2022) achieve efficient and private Transformer inference using Secure Multi-Party Computation (MPC). This measure can mitigate the risk PR1.
5.5. Mitigation on LLM Applications Risks
Secure Auditing and Permission Management. Secure auditing and permission management are crucial for mitigating risks associated with LLM plugins. This approach includes enforcing strict vetting processes to ensure that only trusted plugins are integrated into the system, implementing anomaly detection systems to identify unusual behaviors indicative of malicious activity, and conducting regular security audits to assess plugin security and compliance. Additionally, employing role-based access control (RBAC) limits permissions, ensuring that users and plugins have only the necessary access to perform their functions, thereby reducing the attack surface. By maintaining a comprehensive logging system, organizations can track interactions and quickly respond to potential security incidents, fostering a proactive security posture. This measure can mitigate the risks SR1, SR3 and LR2.
Authorized and Restricted Access. Authorized and restricted access is a key defense mechanism against several risks in LLM applications. For feedback pollution, restricting access to authorized users ensures that only credible feedback is collected for model training. For Middleman (MitM) attacks, by enforcing strong access controls, such as mutual authentication and encryption, only authorized clients can communicate with the LLM. For excessive access, limiting access to verified users via rate-limiting mechanisms mitigates the risk of excessive API requests that can overload the system, leading to Denial of Service (DoS) attacks. By ensuring only authorized users can access the LLM, and controlling the frequency and scope of interactions, systems can prevent abuse and maintain availability. This measure can mitigate the risks DR1, DR2 and DR3.
5.6. Mitigation to Platform Risks
Mitigating Platform Vulnerabilities. Platforms should also ensure that their CI/CD systems are following best security practices, such as least privilege, secure credential storage, and network segmentation, keeping all CI/CD tools and dependencies up to date with the latest security patches to protect against known vulnerabilities (scmagazine, 2024), and regular security assessments and vulnerability scans of the CI/CD environments to promptly identify security flaws. This measure can mitigate the risk SR2.
Session Isolation and Privacy Protection. Using sandboxing in platforms isolates user sessions within separate environments, preventing cross-session attacks. By restricting access to shared resources and data, sandboxing ensures that malicious activities in one session cannot affect others. This containment helps protect sensitive information and enhances overall security by limiting the impact of potential breaches. Besides, platforms can utilize secrets management tools, log scanning, and sanitization to ensure that sensitive information like tokens and API keys are not exposed in repositories or logs (French, 2024). This measure can mitigate the risks SR2, PR1, PR2, PR3 and PR4.
5.7. Mitigation to General LLM Supply Chain Component Risks
@@
Enhancing Security and Legal Risk Awareness. Developers and organizations should be vigilant when incorporating components into their software systems. On the one hand, they should actively monitor for known vulnerabilities in third-party libraries and frameworks, thus reducing their exposure to potential threats. On the other hand, they should familiarize themselves with licensing agreements, intellectual property rights, and the potential liabilities tied to both open-source and proprietary components. This measure can mitigate the risks SR2 and LR1.
@@
Training for Secure Coding. Adhering to secure coding practices is fundamental to minimizing software vulnerabilities. This includes following best practices for input validation, implementing robust error handling, and adhering to established secure coding guidelines. By prioritizing these practices, developers can substantially lower the likelihood of introducing security flaws into their applications, thereby safeguarding user data and assets. This measure can mitigate the risk SR2.
@@
Preparing Alternative Components. To mitigate the risks associated with reliance on components that may become unsupported or discontinued, it is important to seek alternative solutions. Moreover, advanced replacement techniques are also needed to help migrate the original component into a new one with the least effort and cost. By diversifying their dependencies and preparing for potential discontinuations, the stability and security of the LLM-driven application can be enhanced. This measure can mitigate the risks SR2, DR2 and DR3.

### v3

@@ 6.3. Accountability Mechanism
6.4. Resilient Training
@@
6.5. Management
@@
Risk Awareness Training and Secure Auditing. Administrators should be vigilant when components are incorporated into their software systems. On the one hand, they should actively monitor for known vulnerabilities in third-party libraries and frameworks, thus reducing their exposure to potential threats. On the other hand, they should familiarize themselves with licensing agreements, intellectual property rights, and the potential liabilities tied to both open-source and proprietary components. Furthermore, they should conduct regular security audits to assess LLM supply chain component security and compliance. This measure can mitigate the risks SR2 and LR1.
@@
Secure Coding Procedure. Adhering to secure coding practices is fundamental to minimizing software vulnerabilities. This includes following best practices for input validation, implementing robust error handling, and adhering to established secure coding guidelines. By prioritizing these practices, developers can substantially lower the likelihood of introducing security flaws into their applications, thereby safeguarding user data and assets. This measure can mitigate the risk SR2.
Permission Management. Permission management includes enforcing strict vetting processes to ensure that only trusted components are introduced into the system. Additionally, employing role-based access control (RBAC) limits permissions, ensuring that users and LLM supply chain components have only the necessary access to perform their functions, thereby reducing the attack surface. This measure can mitigate the risks SR1, SR3 and LR2.
6.6. Contingency Plan
@@
Preparing Alternatives. To mitigate the risks associated with reliance on components that may become unsupported or discontinued, it is important to seek alternative solutions. Moreover, advanced replacement techniques are also needed to help migrate the original component into a new one with the least effort and cost. By diversifying their dependencies and preparing for potential discontinuations, the stability and security of the LLM-driven application can be enhanced. This measure can mitigate the risks SR2, DR2 and DR3.
Summary. We discusses various strategies to mitigate risks in the LLM supply chain. Key measures include malware detection and Software Bill-of-Materials (SBOM) analysis to address risks from third-party libraries, and privacy-preserving techniques such as data deduplication, encryption, and differential privacy to secure datasets. These strategies address a comprehensive range of risks, ensuring a robust framework for LLM development, deployment, and operation.
7. Case Study
Probllama. Ollama is an open-source and widely-used framework that allows users to operate large language models. Recently, researchers uncovered 6 vulnerabilities in Ollama (oligo, 2024a). The vulnerabilities could allow an attacker to carry out a wide-range of malicious actions, including Denial of Service (DoS) attacks, model poisoning, model stealing, etc. Researchers highlighted that the inherent lack of authentication support in these tools makes them vulnerable when exposed to external environments. The risk scenario is that Model Leakage (PR3), Denial of Service (DR3) are caused by Vulnerability Exploiter stealing models and performing excessive access on AI models, which includes Chatbots, Web Apps and Intelligent Agents, etc. The corresponding mitigation measures include Access Control and Permission Management (oligo, 2024a).
Ray. Ray is a distributed computing framework designed to simplify the process of scaling AI applications. It was reported that attackers have been exploiting a missing authentication vulnerability in the Ray AI framework to compromise hundreds of clusters (oligo, 2024b). By default, Ray does not enforce authentication and does not support any type of authorization methods. Reporters say that hundreds of Ray clusters were hacked via this bug (securityweek, 2024b), with the attackers stealing AI models and data, database credentials, password hashes, SSH keys, and OpenAI, HuggingFace, and Stripe tokens. The risk scenario is that Data Leakage (PR1), Model Leakage (PR3), PII Leakage (PR3) are caused by Vulnerability Exploiter stealing artifacts including data, model and code, etc. The corresponding mitigation measures include Vulnerability Detection and Access Control .
“nullifAI” and Hugging Face Safetensors. Pickle is a popular Python module for serializing and deserializing ML models. However, Pickle is considered an unsafe data format that allows Python code to be executed during deserialization, which can lead to arbitrary code execution. The research team came upon two Hugging Face models containing malicious code exploiting the unsafe mechanisms of Pickle, named as “nullifAI” (reversinglabs, 2024). The risk scenario is that System Compromise (SR2) and Privacy Risks (PR) are caused by Vulnerability Exploiter exploiting artifact vulnerabilities including data, model and code, etc.
Furthermore, the Hugging Face introduces a new safe serialization format called Safetensors (hiddenlayer, 2024) to mitigate the supply chain risk posed by vulnerable serialization formats. They created a conversion service to convert any PyTorch model contained within a repository into a Safetensors alternative via a pull request. However, researchers demonstrate that attackers could compromise the Safetensors conversion space and its associated service bot (hiddenlayer, 2024). The risk scenario is that System Compromise (SR2) and Privacy Risks (PR) are caused by Vulnerability Exploiter exploiting platform vulnerabilities. The corresponding mitigation measures include Vulnerability Detection and Malware Detection.
DeepSeek Brand Impersonation. DeepSeek AI chatbot quickly gained international attention, making it a prime target for abuse. Attackers leverages a tactic known as brand impersonation, creating look-alike websites to deceive users and steal private information (zscaler, 2024; securityweek, 2024a). They also uses a fake CAPTCHA page to deceive users into executing a malicious command. The risk scenario is that System Compromise (SR2) and Privacy Risks (PR) are caused by Vulnerability Exploiter exploiting LLM App. vulnerabilities, and Negligent User introducing Risky Components. The corresponding mitigation measures include Risk Awareness Training and Secure Auditing, Software BOM Analysis and Malware Detection.